Solve
L^-1 × ((s^2 - 5s - 7)/(s + 2)^3)
1
Expert's answer
2020-07-20T18:00:00-0400

Given,

"\mathscr{L^{-1}} \bigg(\dfrac{s^2 - 5s - 7}{(s + 2)^3}\bigg)"

Thus,

"\implies\mathscr{L^{-1}}\left\{\frac{1}{s+2}-\frac{9}{\left(s+2\right)^2}+\frac{7}{\left(s+2\right)^3}\right\}""\implies \mathscr{L^{-1}}\left\{\frac{1}{s+2}\right\}-\mathscr{L^{-1}}\left\{\frac{9}{\left(s+2\right)^2}\right\}+\mathscr{L^{-1}}\left\{\frac{7}{\left(s+2\right)^3}\right\}"

"\implies e^{-2t}-e^{-2t}\cdot \:9t+\frac{7e^{-2t}t^2}{2}"